40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog  Not logged in ELOG logo
Entry  Sun Dec 21 18:46:06 2014, diego, Update, IOO, MC Error Spectra SPAG4395_17-12-2014_170951.pdfSPAG4395_17-12-2014_172846.pdfSPAG4395_17-12-2014_175147.pdfSPAG4395_18-12-2014_003414.pdfSPAG4395_18-12-2014_003506.pdf
    Reply  Mon Dec 22 15:11:08 2014, rana, Update, IOO, MC Error Spectra 
Message ID: 10826     Entry time: Sun Dec 21 18:46:06 2014     Reply to this: 10828
Author: diego 
Type: Update 
Category: IOO 
Subject: MC Error Spectra 

The error spectra I took so far are not that informative, I'm afraid. The first three posted here refer to Wed 17 in the afternoon, where things were quiet, the LSC control was off and the MC was reliably locked. The last two plots refer to Wed night, while Q and I were doing some locking work; in particular, these were taken just after one of the locklosses described in elog 10814. Sadly, they aren't much different from the "quiet" ones.

I can add some considerations though: Q and I saw some weird effects during that night, using a live reading of such spectra, which couldn't be saved though; such effects were quite fast both in appearance and disapperance, therefore difficult to save using the snapshot measurement, which is the only one that can save the data as of now; moreover, these effects were certainly seen during the locklosses, but sometimes also in normal circumstances. What we saw was a broad peak in the range 5e4-1e5 Hz with peak value ~1e-5 V/rtHz, just after the main peak shown in the attached spectra.

Attachment 1: SPAG4395_17-12-2014_170951.pdf  42 kB  | Hide | Hide all | Show all
Attachment 2: SPAG4395_17-12-2014_172846.pdf  42 kB  | Hide | Hide all | Show all
Attachment 3: SPAG4395_17-12-2014_175147.pdf  42 kB  | Hide | Hide all | Show all
Attachment 4: SPAG4395_18-12-2014_003414.pdf  42 kB  | Hide | Hide all | Show all
Attachment 5: SPAG4395_18-12-2014_003506.pdf  42 kB  | Show | Hide all | Show all
ELOG V3.1.3-